# Robust Bayesian Inference for Berkson and Classical Measurement Error Models

### v2

@@ 2 Bayesian Nonparametric Learning
A flexible way of approaching model misspecification in a Bayesian framework is through the Bayesian NPL setting (Lyddon et al.,, 2018; Fong et al.,, 2019). By deriving NPL posteriors, rather than standard Bayesian posteriors, we perform Bayesian inference under misspecified error distributions whilst fully characterizing model uncertainty: prior beliefs about $\operatorname{\mathbb{F}}_{N}^{0}$ are incorporated through a nonparametric prior whereas the model $\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$ is not assumed to be well-specified as in standard Bayesian inference. The resulting framework provides robustness with respect to the parametric distributional specifications, straightforward incorporation of prior beliefs on the ME distribution, and an efficient parallelizable sampling scheme.
NPL relaxes the standard Bayesian assumption that a parametric model family $\operatorname{\mathcal{P}_{\Theta}}=\{\operatorname{\mathbb{P}_{\theta}}:\theta\in\Theta\}$ can correctly describe the DGP $\operatorname{\mathbb{P}}^{0}$, i.e. that $\operatorname{\mathbb{P}}^{0}\in\operatorname{\mathcal{P}_{\Theta}}$. The approach to incorporating uncertainty in the parameter of interest involves placing uncertainty directly on $\operatorname{\mathbb{P}}^{0}$ and propagating it to $\theta\in\Theta$. It is accomplished using a conjugate nonparametric prior and associated posterior on $\operatorname{\mathbb{P}}^{0}$, which then induces a posterior on the parameter space $\Theta$. This differs from traditional Bayesian inference, which conditions on the parameters rather than the DGP. This framework leads to a parallelizable computational framework to sample from this posterior on $\theta$, called the Posterior Bootstrap.
The DGP is modeled using a DP prior, as proposed by Lyddon et al., (2018) and Fong et al., (2019): $\operatorname{\mathbb{P}}\sim\operatorname{\operatorname{DP}}(c,\operatorname{\mathbb{F}})$. Here, $c>0$ is a concentration parameter and $\operatorname{\mathbb{F}}\in\operatorname{\mathcal{P}}$ is a centering measure. Given $(x_{i},y_{i})_{1:n}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{P}}^{0}$, we can obtain the posterior distribution of $\operatorname{\mathbb{P}}$ by conjugacy:
$\begin{split}\operatorname{\mathbb{P}}\>|\>(x_{i},y_{i})_{1:n}\sim\operatorname{\operatorname{DP}}\left(c^{\prime},\operatorname{\mathbb{F}}^{\prime}\right),\quad c^{\prime}=c+n,\qquad\operatorname{\mathbb{F}}^{\prime}=\frac{c}{c+n}\operatorname{\mathbb{F}}+\frac{n}{c+n}\mathbb{P}_{n}\end{split}$ (4)
where $\mathbb{P}_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{(x_{i},y_{i})}$ and $\delta_{(x,y)}$ is a Dirac measure at $(x,y)\in\mathcal{X}\times\mathcal{Y}$. The value of $c$ indicates the level of confidence we have in the prior centering measure $\operatorname{\mathbb{F}}$. The NPL framework, in its standard form, is not sufficient to account for ME, since it assumes that we have access to realizations from the DGP. We introduce an approach, suitable for this setting, based on the NPL principles.
@@ 3 Methodology
We present a Bayesian framework for ME models (MEMs) by placing a nonparametric prior on the conditional distributions $\{X\>|\>W=w_{i}\}_{1:n}$. This is achieved through a flexible DP distribution, allowing the prior definition for both Berkson and Classical error models. We explore two loss functions within this framework: one based on the TLS objective and the other on the MMD. The latter extends the NPL-MMD framework (Dellaporta et al.,, 2022) to the ME context, enabling the handling of non-Gaussian and nonlinear models.
@@ 3.1 NPL for measurement error (Robust-MEM)
The standard NPL setting (Section 2) assumes the availability of a finite number of observations from the DGP and allows for inference on a potentially misspecified model. In the MEM setting, observations from the DGP are not available since we have observations of the pair of random variables $(W,Y)$, instead of $(X,Y)$. In this case, we are interested in setting the uncertainty directly on $X\>|\>W$ and propagating this on $\theta$ through an NPL posterior. Recall the observed data $\{(w_{i},y_{i})\}_{i=1}^{n}$ such that realizations of $X$ are observed with some additive, zero mean, and constant variance error, i.e. $w_{i}=x_{i}+\nu_{i}$ in the Classical case or $x_{i}=w_{i}+\nu_{i}$ in the Berkson case, for all $i=1,\dots,n$. Denote by $\operatorname{\mathbb{P}}_{X\>|\>w_{i}}\in\mathcal{P}(\mathcal{X})$ the distribution of $X\>|\>W=w_{i}$. We set a nonparametric prior on $\operatorname{\mathbb{P}}_{X\>|\>w_{i}}$ through a DP:
@@
The prior centering measure is indexed by $w_{i}$ to indicate the dependence on the observed covariate $w_{i}$. $\operatorname{\mathbb{F}}_{w_{i}}$ encodes our prior beliefs about the true and noisy covariate relationship through distributional assumptions on the ME and the true covariate while respecting the Berkson or Classical structure of the error. We further discuss the choices of $\operatorname{\mathbb{F}}_{w_{i}}$ and $c$ in Sections 3.2 and 3.3 respectively. Since we only have one data point $(w_{i},y_{i})$ for each conditional $X\>|\>W=w_{i}$, the conjugate posterior DP is:
@@
Let $\operatorname{\mathbb{P}}^{0}=\operatorname{\mathbb{P}}_{X}\operatorname{\mathbb{P}}_{Y\>|\>\cdot}\in\mathcal{P}(\mathcal{Z})$ be the true joint DGP where $\operatorname{\mathbb{P}}_{X}\in\mathcal{P}(\mathcal{X})$ denotes the distribution of $X$ and $\operatorname{\mathbb{P}}_{Y\>|\>x}$ denotes the true conditional distribution of $Y\>|\>X=x$ for any $x\in\mathcal{X}$. The target for any loss function $l:\mathcal{X}\times\mathcal{Y}\times\Theta\rightarrow\operatorname{\mathbb{R}}$ is:
@@
Of course, $\operatorname{\mathbb{P}}^{0}$ is unknown and we do not have access to observations from $\operatorname{\mathbb{P}}_{X}$ and consequently $\operatorname{\mathbb{P}}^{0}$. We instead have a set of posterior distributions for each of the conditional distributions $X\>|\>W=w_{i}$. The goal is to use the nonparametric posterior of (6) to obtain a posterior on $\theta\in\Theta$. This is naturally defined through the push-forward measure $(\theta^{*}_{l})^{\#}\left(\mu_{1},\dots,\mu_{n}\right)$ where for probability measures $\operatorname{\mathbb{P}}^{1},\dots,\operatorname{\mathbb{P}}^{n}$ on $\mathcal{P}(\mathcal{X})$:
$\theta^{*}_{l}(\operatorname{\mathbb{P}}^{1},\dots,\operatorname{\mathbb{P}}^{n})$ $:=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{(x,y)\sim\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{P}}^{i}\delta_{y_{i}}}\left[l(x,y,\theta)\right],$ (8)
$\mu_{i}$ $:=\operatorname{\operatorname{DP}}(c+1,\operatorname{\mathbb{F}}_{w_{i}}^{\prime}),\quad\forall i=1,\dots n.$ (9)
Sampling can be performed through the principles of the Posterior Bootstrap (Fong et al.,, 2019) where at iteration j:
@@
    Calculate $\theta^{*}_{l}(\operatorname{\mathbb{P}}^{(1,j)},\dots,\operatorname{\mathbb{P}}^{(n,j)})=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{(x,y)\sim\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{P}}^{(i,j)}\delta_{y_{i}}}\left[l(x,y,\theta)\right]$.
Exact sampling from a DP, as in step 1, is generally not feasible. The stick-breaking process (Sethuraman,, 1994) offers an exact representation of a DP which can be approximated using the Dirichlet approximation of the stick-breaking process (Muliere and Secchi,, 1996; Ishwaran and Zarepour,, 2002) which we use in this work. We denote by $\hat{\mu_{i}}$, the probability measure approximating $\mu_{i}$, meaning that $\operatorname{\mathbb{P}}^{(i,j)}\sim\hat{\mu_{i}}$ is such that:
@@
The intuition of this sampling procedure lies in the fact that different marginal distributions of $\operatorname{\mathbb{P}}^{(i,j)}$ lead to different joint distributions of the form $\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{P}}^{(i,j)}\delta_{y_{i}}$. Uncertainty stems from the marginal distribution of $X$ due to ME, hence by bootstrapping over different marginal distributions of $X$ we induce robustness to the inference procedure for the joint distribution of $(X,Y)$. The taxonomy of the Robust-MEM framework for different choices of loss functions and prior DP specifications is outlined in Figure 2.
@@
3.2 Choice of prior centering measure
The choice of the prior centering measure $\operatorname{\mathbb{F}}_{w_{i}}$ in (5) represents our prior beliefs about the distribution of the true covariate conditionally on an observed noisy covariate $w_{i}$, which should respect the Berkson and Classical dependence assumptions.
@@ 3.2 Choice of prior centering measure
The prior centering measure can then be defined as the probability measure with corresponding density $f_{X\>|\>W}(\cdot\>|\>w_{i})$, which we denote by $\operatorname{\mathbb{F}}_{X\>|\>w_{i}}$. Here, if $f_{X}$ is conjugate to $f_{W\>|\>X}$, then $f_{X\>|\>W}$ can be obtained in closed form and sampling from the corresponding probability measures $\operatorname{\mathbb{F}}_{X\>|\>w_{i}}$ becomes trivial. For example, if $\operatorname{\mathbb{F}}_{X}:=N(\mu_{x},\sigma^{2}_{x})$ and $\operatorname{\mathbb{F}}_{W\>|\>X}:=N(X,\sigma^{2}_{\nu})$ then by conjugacy $\operatorname{\mathbb{F}}_{X\>|\>w_{i}}=N\left(\frac{\sigma_{x}^{2}}{\sigma_{\nu}^{2}+\sigma_{x}^{2}}w_{i}+\frac{\sigma_{\nu}^{2}}{\sigma_{\nu}^{2}+\sigma_{x}^{2}}\mu_{x},(\frac{1}{\sigma_{x}^{2}}+\frac{1}{\sigma_{\nu}^{2}})^{-1}\right)$
Remark 3.2.
The Classical ME structure naturally arises in the study of DP mixture models. Since $X$ is independent of $N$ one could choose to set uncertainty directly to the unknown true distribution of $X$ through a DP and define a mixture model as follows:
$w_{i}\>|\>x_{i}$ $\sim F(x_{i})$
$x_{i}\>|\>G$ $\sim G$
$G$ $\sim\operatorname{\operatorname{DP}}(c,\operatorname{\mathbb{F}})$
where $F$ denotes the distribution of the error conditionally on $x_{i}$ and $G$ is the mixing distribution over $x$. However, such a structure does not lead to a conjugate posterior for $G$ and more complicated sampling procedures should be used such as the ones described in Neal, (2000).
@@
input : $(w_{i},y_{i})_{1:n}$, $T$, $B$, $c$, $l$, $\operatorname{\mathbb{F}}_{N}$
@@
      Obtain $\theta^{(j)}=\theta^{\ast}_{l}(\operatorname{\mathbb{P}}^{(1,j)},\dots,\operatorname{\mathbb{P}}^{(n,j)})$ as in (8) using numerical opt.
return Posterior bootstrap sample $\theta^{(1:B)}$
@@
input : $(w_{i},y_{i})_{1:n}$, $T$, $B$, $c$, $l$, $\operatorname{\mathbb{F}}_{N}$, $\operatorname{\mathbb{F}}_{X}$
for $j\leftarrow 1$ to $B$ do
@@
             Sample $\tilde{x}_{1:T}^{(i,j)}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}_{X\>|\>w_{i}}$
@@
      Obtain $\theta^{(j)}=\theta^{\ast}_{l}(\operatorname{\mathbb{P}}^{(1,j)},\dots,\operatorname{\mathbb{P}}^{(n,j)})$ as in (8) using numerical opt.
return Posterior bootstrap sample $\theta^{(1:B)}$

### v3

@@ 2 Bayesian Nonparametric Learning
A flexible way of approaching model misspecification in a Bayesian framework is through the Bayesian NPL setting (Lyddon et al.,, 2018; Fong et al.,, 2019). By deriving NPL, rather than standard Bayesian, posteriors we perform Bayesian inference under misspecified error distributions whilst fully characterising model uncertainty: prior beliefs about X|WX\>|\>W are incorporated through a nonparametric prior. The resulting framework offers robustness against parametric distributional assumptions and facilitates the inclusion of prior beliefs regarding the error distribution.
NPL relaxes the standard Bayesian assumption that a parametric model family $\operatorname{\mathcal{P}_{\Theta}}$ can correctly describe the Data-Generating Process (DGP) $\operatorname{\mathbb{P}}^{0}$, i.e. that $\operatorname{\mathbb{P}}^{0}\in\operatorname{\mathcal{P}_{\Theta}}$. The approach to incorporating uncertainty in the parameter of interest involves placing uncertainty directly on $\operatorname{\mathbb{P}}^{0}$ and propagating it to $\theta\in\Theta$. It is accomplished using a conjugate nonparametric prior and associated posterior on $\operatorname{\mathbb{P}}^{0}$, which then induces a posterior on the parameter space $\Theta$. This differs from traditional Bayesian inference, which conditions on the parameters rather than the DGP. This framework leads to a parallelizable computational framework to sample from this posterior on $\theta$, called the Posterior Bootstrap.
[…]
$\begin{split}\textstyle\operatorname{\mathbb{P}}\>|\>\{(x_{i},y_{i})\}_{i=1}^{n}\sim\operatorname{\operatorname{DP}}\left(c^{\prime},\operatorname{\mathbb{F}}^{\prime}\right),\quad c^{\prime}=c+n,\qquad\operatorname{\mathbb{F}}^{\prime}=\frac{c}{c+n}\operatorname{\mathbb{F}}+\frac{n}{c+n}\mathbb{P}_{n}\end{split}$ (2)
where $\mathbb{P}_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{(x_{i},y_{i})}$ and $\delta_{(x,y)}$ is a Dirac measure at $(x,y)\in\mathcal{Z}$. The value of $c$ indicates the level of confidence we have in the prior centring measure $\operatorname{\mathbb{F}}$. The NPL framework, in its standard form, is not sufficient to account for ME, since it assumes that we have access to realisations from the DGP. We introduce an approach, suitable for this setting, based on the NPL principles.
@@ 3 Methodology
We present a Bayesian framework for ME models (MEMs) by means of a nonparametric prior on the conditional distributions $\{X\>|\>W=w_{i}\}_{i=1}^{n}$. This is achieved through a flexible DP distribution, allowing the prior definition for both Berkson and Classical error models. We explore two loss functions within this framework: one based on the TLS objective and the other on the MMD. The latter extends the NPL-MMD framework (Dellaporta et al.,, 2022) to the ME context, enabling the handling of non-Gaussian and nonlinear models.
@@ 3.1 NPL for Measurement Error (Robust-MEM)
The standard NPL setting (Sec. 2) assumes the availability of a finite number of observations from the DGP and allows for inference on a potentially misspecified model. In the MEM setting, observations from the DGP are not available since we have observations of the pair of random variables $(W,Y)$, instead of $(X,Y)$. In this case, we are interested in setting the uncertainty directly on X|WX\>|\>W and propagating this on $\theta$ through an NPL posterior. Recall the observed data $\{(w_{i},y_{i})\}_{i=1}^{n}$, such that realizations of $X$ are observed with some additive, zero mean, and constant variance error, i.e. $w_{i}=x_{i}+\nu_{i}$ in the Classical case or $x_{i}=w_{i}+\nu_{i}$ in the Berkson case, for all $i=1,\dots,n$. Denote by $\operatorname{\mathbb{P}}_{X\>|\>w_{i}}\in\mathcal{P}_{\mathcal{X}}$ the distribution of $X\>|\>W=w_{i}$. We set a nonparametric prior on $\operatorname{\mathbb{P}}_{X\>|\>w_{i}}$ through a DP:
@@
The prior centering measure $\operatorname{\mathbb{F}}_{w_{i}}$ is indexed by $w_{i}$ to indicate the dependence on the observed covariate $w_{i}$. $\operatorname{\mathbb{F}}_{w_{i}}$ encodes our prior beliefs about the true and noisy covariate relationship through distributional assumptions on the ME and the true covariate while respecting the Berkson or Classical structure of the error. We further discuss the choices of $\operatorname{\mathbb{F}}_{w_{i}}$ and $c$ in Sections 3.2 and 3.3 respectively. Since we only have one data point $(w_{i},y_{i})$ for each conditional $X\>|\>W=w_{i}$, the conjugate posterior DP is:
@@
Let $\operatorname{\mathbb{P}}^{0}=\operatorname{\mathbb{P}}^{0}_{X}\times\operatorname{\mathbb{P}}^{0}_{Y\>|\>X}\in\mathcal{P}_{\mathcal{Z}}$ be the true joint DGP where $\operatorname{\mathbb{P}}^{0}_{X}\in\mathcal{P}_{\mathcal{X}}$ denotes the distribution of $X$ and $\operatorname{\mathbb{P}}^{0}_{Y\>|\>X}$ denotes the true conditional distribution of Y|XY\>|\>X. The target for any loss function $l:\mathcal{X}\times\mathcal{Y}\times\Theta\rightarrow\operatorname{\mathbb{R}}$ is:
@@
Of course, $\operatorname{\mathbb{P}}^{0}$ is unknown and we do not have access to observations from $\operatorname{\mathbb{P}}_{X}^{0}$ and consequently $\operatorname{\mathbb{P}}^{0}$. We instead have a set of posterior distributions for each of the conditional distributions $X\>|\>W=w_{i}$. The goal is to use the nonparametric posterior of (4) to obtain a posterior on $\theta\in\Theta$. This is naturally defined through the pushforward measure:
$(\theta^{\star}_{l})^{\#}\left(\mu_{1},\dots,\mu_{n}\right)$
where for probability measures $\operatorname{\mathbb{P}}^{1}\sim\mu_{1},\dots,\operatorname{\mathbb{P}}^{n}\sim\mu_{n}$ on $\mathcal{P}_{\mathcal{X}}$ and $\theta^{\star}_{l}:\mathcal{P}_{\mathcal{X}}^{n}\rightarrow\Theta$ we define:
$\operatorname{\mathbb{P}}^{i}\sim\mu_{i}$ $:=\operatorname{\operatorname{DP}}(c+1,\operatorname{\mathbb{F}}_{w_{i}}^{\prime})\quad\forall i=1,\dots n,$ (6)
$\theta^{\star}_{l}(\operatorname{\mathbb{P}}^{1},\dots,\operatorname{\mathbb{P}}^{n})$ $:=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{(x,y)\sim\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{P}}^{i}\times}\delta_{y_{i}}\left[l(x,y,\theta)\right].$ (7)
Sampling from $(\theta^{\star}_{l})^{\#}\left(\mu_{1},\dots,\mu_{n}\right)$ can be performed through the principles of the Posterior Bootstrap (Fong et al.,, 2019) where at iteration j:
@@
    $\theta^{\star}_{l}(\operatorname{\mathbb{P}}^{(1,j)},\dots,\operatorname{\mathbb{P}}^{(n,j)})=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{(x,y)\sim\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{P}}^{(i,j)}\times}\delta_{y_{i}}\left[l(x,y,\theta)\right]$.
Exact sampling from a DP, as in step 1, is generally not feasible. The stick-breaking process (Sethuraman,, 1994) offers an exact representation of a DP which can be approximated using the Dirichlet approximation of the stick-breaking process (Muliere and Secchi,, 1996; Ishwaran and Zarepour,, 2002) which we use in this work. We denote by $\hat{\mu_{i}}$, the approximation of $\mu_{i}$, meaning that $\operatorname{\mathbb{P}}^{(i,j)}\sim\hat{\mu_{i}}$ is such that:
@@
The intuition of this sampling procedure lies in the fact that different marginal distributions of $\operatorname{\mathbb{P}}^{(i,j)}$ lead to different joint distributions of the form
$\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{P}}^{(i,j)}\times\delta_{y_{i}}.$
Uncertainty stems from the marginal distribution of $X$ due to ME, hence by bootstrapping over different marginal distributions of $X$ we induce robustness to the inference procedure for the joint distribution of $(X,Y)$. The taxonomy of the Robust-MEM framework for different choices of loss functions and prior DP specifications is outlined in Fig. 2.
3.2 Choice of Prior Centering Measure
The choice of the prior centering measure $\operatorname{\mathbb{F}}_{w_{i}}$ in (3) represents our prior beliefs about the distribution of the true covariate conditionally on an observed noisy covariate $w_{i}$. This choice should respect the Berkson and Classical dependence assumptions.
@@ 3.2 Choice of Prior Centering Measure
$f_{X\>|\>W}(x\>|\>w_{i})=\frac{f_{W\>|\>X}(w_{i}\>|\>x)f_{X}(x)}{\int f_{W\>|\>X}(w_{i}\>|\>x)f_{X}(x)dx}$ (9)
The prior centring measure can then be defined as the probability measure with corresponding density $f_{X\>|\>W}(\cdot\>|\>w_{i})$, which we denote by $\operatorname{\mathbb{F}}_{X\>|\>w_{i}}$. Here, if $f_{X}$ is conjugate to $f_{W\>|\>X}$, then $f_{X\>|\>W}$ can be obtained in closed form and sampling from the corresponding probability measures $\operatorname{\mathbb{F}}_{X\>|\>w_{i}}$ becomes trivial. For example, if $\operatorname{\mathbb{F}}_{X}:=\mathcal{N}(\mu_{x},\sigma^{2}_{x})$ and $\operatorname{\mathbb{F}}_{W\>|\>X}:=\mathcal{N}(X,\sigma^{2}_{\nu})$ then by conjugacy $\operatorname{\mathbb{F}}_{X\>|\>w_{i}}=\mathcal{N}\left(\frac{\sigma_{x}^{2}}{\sigma_{\nu}^{2}+\sigma_{x}^{2}}w_{i}+\frac{\sigma_{\nu}^{2}}{\sigma_{\nu}^{2}+\sigma_{x}^{2}}\mu_{x},(\frac{1}{\sigma_{x}^{2}}+\frac{1}{\sigma_{\nu}^{2}})^{-1}\right)$.
Remark 3.2.
@@
$w_{i}\>|\>x_{i}\sim F(x_{i}),\quad x_{i}\>|\>G$ $\sim G,\quad G\sim\operatorname{\operatorname{DP}}(c,\operatorname{\mathbb{F}}),$
where $F$ denotes the distribution of the error conditionally on $x_{i}$ and $G$ is the mixing distribution over $x$. However, such a structure does not lead to a conjugate posterior for $G$ and more complicated sampling procedures should be used (Neal,, 2000).
@@
input : $\{(w_{i},y_{i})\}_{i=1}^{n}$, $T$, $B$, $c$, $l$, $\operatorname{\mathbb{F}}_{N}$
@@
 $\theta^{(j)}=\theta^{\ast}_{l}(\operatorname{\mathbb{P}}^{(1,j)},\dots,\operatorname{\mathbb{P}}^{(n,j)})$ in (7)
return Posterior sample $\theta^{(1:B)}$
@@
input : $\{(w_{i},y_{i})\}_{i=1}^{n}$, $T$, $B$, $c$, $l$, $\operatorname{\mathbb{F}}_{N}$, $\operatorname{\mathbb{F}}_{X}$
0.07cm for $j\leftarrow 1$ to $B$ do
@@
      Set $\tilde{x}_{1:T}^{(i,j)}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}_{X\>|\>w_{i}}$
@@
 $\theta^{(j)}=\theta^{\ast}_{l}(\operatorname{\mathbb{P}}^{(1,j)},\dots,\operatorname{\mathbb{P}}^{(n,j)})$ in (7)
return Posterior sample $\theta^{(1:B)}$